OpenStudy (anonymous):

Type the equation of the given line in standard form. The line with a slope of (-1/3) and b of 6.

OpenStudy (anonymous):

y=(-1/3)x+6

OpenStudy (anonymous):

y=mx+b....... where m is the slope and b is y-intercept

OpenStudy (anonymous):

And then to put it into standard form I just rearrange everything?

OpenStudy (phi):

Of course, you have to make it into standard form http://en.wikipedia.org/wiki/Linear_equation#Standard_form

OpenStudy (anonymous):

y=(-1/3)x + 6 turns into (-1/3)x + y = 6 Right?

OpenStudy (phi):

you add +1/3 x to both sides this makes it y + 1/3 x = -1/3 x + 1/3 x +6 (and the right side simplifies) 1/3 x + y = 6 (standard form has the x term first)

OpenStudy (anonymous):

Wait. No, it would be +(1/3).

OpenStudy (anonymous):

Whoops, you just said that :P

OpenStudy (anonymous):

Thank you!